# Multi-Armed Bandits with Abstention

### 23 Feb 2024 v1

@@ Abstract
We introduce a novel extension of the canonical multi-armed bandit problem that incorporates an additional strategic element: abstention. In this enhanced framework, the agent is not only tasked with selecting an arm at each time step, but also has the option to abstain from accepting the stochastic instantaneous reward before observing it. When opting for abstention, the agent either suffers a fixed regret or gains a guaranteed reward. Given this added layer of complexity, we ask whether we can develop efficient algorithms that are both asymptotically and minimax optimal. We answer this question affirmatively by designing and analyzing algorithms whose regrets meet their corresponding information-theoretic lower bounds. Our results offer valuable quantitative insights into the benefits of the abstention option, laying the groundwork for further exploration in other online decision-making problems with such an option. Numerical results further corroborate our theoretical findings.
Machine Learning, ICML
@@ 1 Introduction
In the realm of online decision-making, the multi-armed bandit model, originally introduced by Thompson (1933), has long served as a quintessential benchmark for capturing the delicate interplay between exploration and exploitation. In stochastic multi-armed bandit problems, the agent sequentially selects an arm from the given set at each time step and subsequently observes a random reward associated with the chosen arm. To maximize cumulative rewards, the agent must strike a balance between the persistent pursuit of the arm with the highest expected reward (exploitation) and the adventurous exploration of other arms to gain a deeper understanding of their potential (exploration). This fundamental challenge finds applications across a wide array of domains, ranging from optimizing advertising campaigns to fine-tuning recommendation systems.
@@
However, real-world scenarios often come fraught with complexities that challenge the simplicity of the canonical bandit model. One notable complexity arises when the agent is equipped with an additional option to abstain from accepting the stochastic instantaneous reward before actually observing it. This added layer of decision-making considerably enriches the strategic landscape, altering how the agent optimally navigates the trade-off between exploration and exploitation.
@@
Consider, for example, the domain of clinical trials. When evaluating potentially hazardous medical treatments, researchers can proactively deploy safeguards such as preemptive medications or consider purchasing specialized insurance packages to shield against possible negative consequences. However, these protective measures come with costs, which may be modeled as either fixed regrets or fixed rewards in the context of the clinical study’s cumulative regret. In these scenarios, researchers have the option to observe the outcomes of a treatment while abstaining from incurring the associated random regret through these costly prearranged measures. Subsequently, these observed outcomes can be leveraged to inform more efficacious future experimental designs. Consequently, opting for abstention has the potential to promote more responsible decision-making and reduce the overall cumulative regret of the study.
@@
    In Section 2, we provide a rigorous mathematical formulation of the multi-armed bandit model with abstention. Our focus is on cumulative regret minimization across two distinct yet complementary settings: fixed-regret and fixed-reward. These settings give rise to divergent performance metrics, each offering unique analytical insights. Importantly, both settings encompass the canonical bandit model as a particular case.
@@
    In the fixed-regret setting, we judiciously integrate two abstention criteria into a Thompson Sampling-based algorithm proposed by Jin et al. (2023). This integration ensures compatibility with the abstention option, as elaborated in Algorithm 1. The first abstention criterion employs a carefully constructed lower confidence bound, while the second is tailored to mitigate worst-case scenarios. We establish both asymptotic and minimax upper bounds on the cumulative regret. Furthermore, we derive corresponding lower bounds, thereby demonstrating that our algorithm attains asymptotic and minimax optimality simultaneously.
@@
    In the fixed-reward setting, we introduce a general strategy, outlined in Algorithm 2. This method is capable of transforming any algorithm that is both asymptotically and minimax optimal in the canonical model to one that also accommodates the abstention option. Remarkably, this strategy maintains its universal applicability and straightforward implementation while provably achieving both forms of optimality—asymptotic and minimax.
@@
    To empirically corroborate our theoretical contributions, we conduct a series of numerical experiments in Section 5. These experiments substantiate the effectiveness of our algorithms and highlight the performance gains achieved through the inclusion of the abstention option.
@@ 1.1 Related Work
[…]
@@
Starting with the seminal works of Chow (1957, 1970), the concept of learning with abstention (also referred to as rejection) has been extensively explored in various machine learning paradigms. These include, but are not limited to, classification (Herbei & Wegkamp, 2006; Bartlett & Wegkamp, 2008; Cortes et al., 2016), ranking (Cheng et al., 2010; Mao et al., 2023), and regression (Wiener & El-Yaniv, 2012; Zaoui et al., 2020; Kalai & Kanade, 2021).
@@
Although set in different contexts, these existing works consistently demonstrate the value of incorporating abstention into online decision-making processes, underscoring the urgent need to analyze and quantify its benefits in the field of multi-armed bandits.
@@ 2 Problem Setup
We consider a $K$-armed bandit model, enhanced with an additional option to abstain from accepting the stochastic instantaneous reward prior to its observation. Let $\mu\in\mathcal{U}:=\mathbb{R}^{K}$ denote a specific bandit instance, where $\mu_{i}$ represents the unknown mean reward associated with pulling arm $i\in[K]$. For simplicity, we assume that arm $1$ is the unique optimal arm, i.e., $1=\operatorname*{arg\,max}_{i\in[K]}\mu_{i}$, and we define $\Delta_{i}:=\mu_{1}-\mu_{i}$ as the suboptimality gap for each arm $i\in[K]$.
At each time step $t\in\mathbb{N}$, the agent chooses an arm $A_{t}$ from the given arm set $[K]$, and, simultaneously, decides whether or not to abstain, indicated by a binary variable $B_{t}$. Regardless of the decision to abstain, the agent observes a random variable $X_{t}$ from the selected arm $A_{t}$, which is drawn from a Gaussian distribution $\mathcal{N}(\mu_{A_{t}},1)$ and independent of observations obtained from the previous time steps. Notably, the selection of both $A_{t}$ and $B_{t}$ might depend on the previous decisions and observations, as well as on each other. More formally, let $\mathcal{F}_{t}:=\sigma(A_{1},B_{1},X_{1},\ldots,A_{t},B_{t},X_{t})$ denote the $\sigma$-field generated by the cumulative interaction history up to and including time $t$. It follows that the pair of random variables $(A_{t},B_{t})$ is $\mathcal{F}_{t-1}$-measurable.
@@
Alternatively, in the fixed-reward setting, the reward of the abstention option is predetermined to be $c\in\mathbb{R}$ [^1] With a slight abuse of notation, we employ the symbol $c$ to represent the abstention regret and the abstention reward within their respective settings. The surrounding context, nonetheless, should elucidate the exact meaning of $c$. Since the abstention reward $c$ may potentially surpass $\mu_{1}$, the best possible expected reward at a single time step is $\mu_{1}\vee c:=\max\{\mu_{1},c\}$. If the agent decides to abstain ($B_{t}=1$), it guarantees a deterministic reward of $c$, leading to a regret of $\mu_{1}\vee c-c$. Conversely, if $B_{t}=0$, the agent receives a per-time reward $X_{t}$, resulting in a regret of $\mu_{1}\vee c-X_{t}$.
ptAgentptActionptEnvironmentptInstantaneousRegretpt
@@
Remark 2.1.
Regarding the learnability of the abstention model, if the agent cannot observe the sample from the chosen arm when opting for abstention, then it is equivalent to skipping the time step, rendering the model trivial. In other words, the learner gains no information from the time step but incurs an instantaneous regret. In line with this model assumption, our model well aligns with that of Neu & Zhivotovskiy (2020), which explored the role of abstention in the context of online prediction with expert advice.
@@
Our overarching goal is to design and analyze online algorithms $\pi$ that minimize their expected cumulative regrets up to and including the time horizon $T$ [^2] In certain real-world applications, the time horizon $T$ may be unknown to the agent. In fact, all of our proposed methods are inherently anytime in nature, as they do not necessitate prior knowledge of the horizon. The regrets are formally defined for the two distinct settings as follows:
@@
$R^{\mathrm{RG}}_{\mu,c}(T,\pi):=\mathbb{E}\Bigg{[}$ $\sum_{t=1}^{T}\big{(}(\mu_{1}-X_{t})\cdot\mathbbm{1}\{B_{t}=0\}$ (1)
$+c\cdot\mathbbm{1}\{B_{t}=1\}\big{)}\Bigg{]}.$
@@
$\!\!\!R^{\mathrm{RW}}_{\mu,c}(T,\pi):=T\!\cdot\!\left(\mu_{1}\vee c\right)-\mathbb{E}\Bigg{[}$ $\sum_{t=1}^{T}\Big{(}X_{t}\cdot\mathbbm{1}\{B_{t}=0\}\!$ (2)
$+c\cdot\mathbbm{1}\{B_{t}=1\}\Big{)}\Bigg{]}.$
@@
Remark 2.2.
It is worth mentioning that our model is a strict generalization of the canonical multi-armed bandit model (without the abstention option). Specifically, it particularizes to the canonical model as the abstention regret $c$ tends to positive infinity in the fixed-regret setting and as the abstention reward $c$ tends to negative infinity in the fixed-reward setting. Nevertheless, the incorporation of an extra challenge, the abstention decision (denoted as $B_{t}$), offers the agent the potential opportunity to achieve superior performance in terms of either regret.
@@
For $x,y\in\mathbb{R}$, we denote $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. For any arm $i\in[K]$, let $N_{i}(t):=\sum_{s=1}^{t}\mathbbm{1}\{A_{s}=i\}$ and $\hat{\mu}_{i}(t):=\sum_{s=1}^{t}X_{s}\mathbbm{1}\{A_{s}=i\}/N_{i}(t)$ denote its total number of pulls and empirical estimate of the mean up to time $t$, respectively. In particular, we set $\hat{\mu}_{i}(t)=+\infty$ if $N_{i}(t)=0$. To count abstention records, we also use $N_{i}^{(0)}(t)$ and $N_{i}^{(1)}(t)$ to denote its number of pulls without and with abstention up to time $t$, respectively. That is, $N_{i}^{(0)}(t):=\sum_{s=1}^{t}\mathbbm{1}\{A_{s}=i\text{ and }B_{s}=0\}$ and $N_{i}^{(1)}(t):=\sum_{s=1}^{t}\mathbbm{1}\{A_{s}=i\text{ and }B_{s}=1\}$. Additionally, we define $\hat{\mu}_{is}$ as the empirical mean of arm $i$ based on its first $s$ pulls. Furthermore, we use $\alpha$, $\alpha_{1}$, and so forth to represent universal constants that do not depend on the problem instances (including $\mu$, $c$, $T$, $K$), with possibly different values in different contexts.
@@ 3 Fixed-Regret Setting
In terms of the arm sampling rule, our algorithm is built upon Less-Exploring Thompson Sampling (Jin et al., 2023), a minimax optimal enhancement of the celebrated Thompson Sampling algorithm (Thompson, 1933). We refer to Remark 3.1 for the reason behind this choice. During the initialization phase, each arm is sampled exactly once. Following that, at each time $t$, an estimated reward $a_{i}(t)$ is constructed for each arm $i\in[K]$, which is either drawn from the posterior distribution $\mathcal{N}(\hat{\mu}_{i}(t-1),1/N_{i}(t-1))$ with probability $1/K$ or set to be the empirical mean $\hat{\mu}_{i}(t-1)$ otherwise. Subsequently, the algorithm consistently pulls the arm $A_{t}$ with the highest estimated reward.
With regard to the abstention decision rule, we propose two abstention criteria that work in tandem (as detailed in Step 5 of Algorithm 1). The first criterion is gap-dependent in nature. In particular, for each arm $i\in[K]$, we define its lower confidence bound as
@@
Then we choose to abstain if there exists an arm $i\in[K]\setminus\{A_{t}\}$ for which the difference between $\mathrm{LCB}_{i}(t)$ and the empirical mean of the arm $A_{t}$ exceeds $c$. This condition signifies that the suboptimality gap $\Delta_{A_{t}}$ is at least $c$ with high probability. The second abstention criterion is gap-independent and more straightforward. It is motivated from the construction of worst-case scenarios as detailed in the proof of our lower bound. Under this criterion, we opt for the abstention option if $c\leq\sqrt{K/t}$, which implies that the abstention regret remains acceptably low at time $t$ in view of the worst-case scenarios.
Remark 3.1.
As previously highlighted, our model in the fixed-regret setting particularizes to the canonical multi-armed bandit model as the abstention regret $c$ approaches infinity. Similarly, when $c$ tends towards infinity, the two abstention criteria are never satisfied, and the procedure of Algorithm 1 simplifies to that of Less-Exploring Thompson Sampling. It is worth noting that this latter algorithm is not only asymptotically optimal but also minimax optimal for the canonical model. This is precisely why we base our algorithm upon it, rather than the conventional Thompson Sampling algorithm, which has been shown not to be minimax optimal (Agrawal & Goyal, 2017).
Remark 3.2.
To preserve the optimality properties of Less-Exploring Thompson Sampling, it is required to carefully manage the simultaneous control of asymptotic and worst-case regrets resulting from incorrectly opting for the abstention option. The lower confidence bound defined in Equation (3) is meticulously tailored for our specific application. Notably, our choice to employ this LCB rather than other natural types of upper confidence bounds in the abstention criteria is primarily driven by the necessity to control the magnitude of worst-case regret.

### 22 Mar 2026 v2

# Asymptotically and Minimax Optimal Regret Bounds for Multi-Armed Bandits with Abstention

@@ Abstract
We introduce a novel extension of the canonical multi-armed bandit problem that incorporates an additional strategic innovation: abstention. In this enhanced framework, the agent is not only tasked with selecting an arm at each time step, but also has the option to abstain from accepting the stochastic instantaneous reward before observing it. When opting for abstention, the agent either suffers a fixed regret or gains a guaranteed reward. This added layer of complexity naturally prompts the key question: can we develop algorithms that are both computationally efficient and asymptotically and minimax optimal in this setting? We answer this question in the affirmative by designing and analyzing algorithms whose regrets meet their corresponding information-theoretic lower bounds. Our results offer valuable quantitative insights into the benefits of the abstention option, laying the groundwork for further exploration in other online decision-making problems with such an option. Extensive numerical experiments validate our theoretical results, demonstrating that our approach not only advances theory but also has the potential to deliver significant practical benefits.
@@ 1 Introduction
In online decision-making, the multi-armed bandit model, originally introduced by Thompson (1933), has long served as a quintessential benchmark for capturing the delicate interplay between exploration and exploitation. In stochastic multi-armed bandit problems, the agent sequentially selects an arm from the given set at each time step and subsequently observes a random reward associated with the chosen arm. To maximize cumulative rewards, the agent must strike a balance between the persistent pursuit of the arm with the highest estimated reward (exploitation) and the adventurous exploration of other arms to gain a deeper understanding of their potential (exploration). This fundamental challenge finds applications across a wide array of domains, ranging from optimizing advertising campaigns to fine-tuning recommendation systems.
@@
However, real-world scenarios often come fraught with complexities that challenge the simplicity of the canonical bandit model. One notable complexity arises when the agent is equipped with an additional option to abstain from accepting the stochastic instantaneous reward before observing it. This added layer of decision-making enriches the strategic landscape, altering how the agent optimally navigates the trade-off between exploration and exploitation.
@@
Consider, for example, the domain of clinical trials. [^1] We highlight that this example is a stylized mathematical abstraction used solely to motivate the abstention framework. It does not capture the complex ethical and practical constraints of actual clinical trials and should not be interpreted as a proposal for real-world medical protocols. When evaluating potentially hazardous medical treatments, researchers can proactively deploy safeguards such as preemptive medications or consider purchasing specialized insurance packages to shield against possible negative consequences. However, these protective measures come with costs, which may be modeled as either fixed regrets or fixed rewards in the context of the clinical study’s cumulative regret. In these scenarios, researchers have the option to observe the outcomes of a treatment while abstaining from incurring the associated random regret through these costly prearranged measures. Subsequently, these observed outcomes can be leveraged to inform more efficacious future experimental designs. Consequently, opting for abstention has the potential to promote more responsible decision-making and reduce the overall cumulative regret of the study.
To further illustrate our motivation, consider another concrete example in the realm of online advertising, where companies frequently grapple with the challenge of allocating their advertising budget optimally across multiple platforms (Google Ads, LinkedIn, etc.) to promote their products. This scenario naturally parallels the framework of multi-armed bandits, with each arm representing a distinct advertising platform. Typically, these platforms operate on a cost-per-click pricing model, where payment is made for each click generated. However, what companies truly care about is the outcome, such as conversions or purchases, which are random variables sampled from unknown distributions. When faced with such uncertainty and the possibility of underperformance, companies may instead opt for a cost-per-action strategy, which involves a predetermined payment for a specific outcome (i.e., payment is made only when a desired outcome, such as a conversion or purchase, occurs). This strategy essentially serves as an abstention option, where instead of risking uncertain returns (from unknown distributions), companies choose to pay a fixed cost for a known outcome. By embracing the abstention option, companies can mitigate the uncertainties associated with stochastic outcomes and ensure a deterministic result. Moreover, by opting for abstention, the company still retains the ability to observe the advertising conversion rate of the platform, enabling them to gather valuable insights for future decision-making.
@@
    In Section 2, we provide a rigorous formulation of the multi-armed bandit model with abstention. Our focus is on cumulative regret minimization across two distinct yet complementary settings: fixed-regret and fixed-reward. These settings give rise to different metrics, each offering unique analytical insights. Importantly, both settings encompass the canonical bandit model as a particular case.
@@
    In the fixed-regret setting, we judiciously combine two abstention criteria into a Thompson Sampling-based algorithm proposed by Jin et al. (2023). This integration ensures compatibility with the abstention option, as elaborated in Algorithm 1. The first abstention criterion employs a carefully constructed lower confidence bound, while the second is tailored to mitigate worst-case scenarios. We establish both asymptotic and minimax upper bounds on the cumulative regret, requiring application of analytical techniques inspired by both Thompson Sampling and UCB methodologies. Furthermore, we derive corresponding lower bounds, thereby showcasing the concurrent attainment of asymptotic and minimax optimality by our algorithm. Our approach entails extensive application of analytical techniques inspired by both Thompson Sampling and UCB methodologies.
@@
    In the fixed-reward setting, we introduce a general strategy, outlined in Algorithm 2. This method transforms any algorithm that is both asymptotically and minimax optimal in the canonical model to one that also accommodates the abstention option. Remarkably, this strategy maintains its universal applicability and straightforward implementation while provably achieving both forms of optimality—asymptotic and minimax.
@@
    To corroborate our theoretical contributions, we conduct a series of numerical experiments in Appendix F. These experiments substantiate the effectiveness of our algorithms and highlight the performance gains achieved through the inclusion of the abstention option.
@@ 1.1 Related Work
The study of cumulative regret minimization in canonical multi-armed bandits has attracted considerable scholarly focus. Within this domain, two dominant paradigms for evaluating optimality metrics have emerged: asymptotic optimality and minimax optimality. In essence, the former considers the behavior of algorithms as the time horizon approaches infinity for a specific problem instance, while the latter seeks to minimize the worst-case regret over all possible instances. A diverse array of policies have been rigorously established to achieve asymptotic optimality across various model settings. Notable examples include UCB2 (Auer et al., 2002), DMED (Honda & Takemura, 2010), KL-UCB (Cappé et al., 2013), and Thompson Sampling (Agrawal & Goyal, 2012; Kaufmann et al., 2012). In the context of the worst-case regret, MOSS (Audibert & Bubeck, 2009) stands out as the pioneering method that has been verified to be minimax optimal. Remarkably, $\text{KL-UCB}^{++}$ (Ménard & Garivier, 2017) became the first algorithm proved to achieve both asymptotic and minimax optimality. Recently, Jin et al. (2023) introduced Less-Exploring Thompson Sampling, an innovation that boosts computational efficiency compared to classical Thompson Sampling while concurrently achieving asymptotic and minimax optimality. For a comprehensive survey of bandit algorithms, we refer to Lattimore & Szepesvári (2020).
@@
Starting with the seminal work of Chow (1957; 1970), the concept of learning with abstention (also referred to as rejection) has been extensively explored in various machine learning paradigms. These include classification (Herbei & Wegkamp, 2006; Bartlett & Wegkamp, 2008; Cortes et al., 2016), ranking (Cheng et al., 2010; Mao et al., 2023) and regression (Wiener & El-Yaniv, 2012; Zaoui et al., 2020; Kalai & Kanade, 2021).
@@ 2 Problem Setup
We consider a $K$-armed bandit model, enhanced with an additional option to abstain from accepting the stochastic instantaneous reward prior to its observation. Let $\mu\in\mathcal{U}:=\mathbb{R}^{K}$ denote a specific bandit instance, where $\mu_{i}$ represents the unknown mean reward of arm $i\in[K]$. We assume that arm $1$ is the unique optimal arm, i.e., $1=\operatorname*{arg\,max}_{i\in[K]}\mu_{i}$, and we define $\Delta_{i}:=\mu_{1}-\mu_{i}$ as the suboptimality gap for each arm $i$.
At each time step $t\in\mathbb{N}$, the agent chooses an arm $A_{t}\in[K]$, and, simultaneously, decides whether or not to abstain, indicated by a binary variable $B_{t}$. Regardless of the decision to abstain, the agent observes a random variable $X_{t}$ from the selected arm $A_{t}$, which is drawn from a Gaussian distribution $\mathcal{N}(\mu_{A_{t}},1)$ independent of observations obtained from the previous time steps. Notably, the selection of both $A_{t}$ and $B_{t}$ might depend on the previous decisions and observations, as well as on each other. Formally, let $\mathcal{F}_{t}:=\sigma(A_{1},B_{1},X_{1},\ldots,A_{t},B_{t},X_{t})$ denote the $\sigma$-field generated by the cumulative interaction history up to and including time $t$. It follows that the pair of random variables $(A_{t},B_{t})$ is $\mathcal{F}_{t-1}$-measurable.
@@
In the fixed-reward setting, the reward of the abstention option is $c\in\mathbb{R}$. [^2] With a slight abuse of notation, we use the symbol $c$ to represent the abstention regret and the abstention reward within their respective settings. The surrounding context should elucidate the exact meaning of $c$. Since the abstention reward $c$ may be larger than $\mu_{1}$, the highest expected reward at each time step is $\mu_{1}\vee c:=\max\{\mu_{1},c\}$. If the agent decides to abstain ($B_{t}=1$), it is guaranteed a deterministic reward of $c$, leading to a regret of $\mu_{1}\vee c-c$. If $B_{t}=0$, the agent receives a reward of $X_{t}$, resulting in a regret of $\mu_{1}\vee c-X_{t}$.
AgentActionEnvironmentInstantaneousRegret
@@
Remark 1.
We note that if the agent cannot observe the sample from the chosen arm when opting for abstention, then it is equivalent to skipping the time step, rendering the model trivial. In other words, the learner gains no information from the time step but incurs an instantaneous regret. In view of this, our model well aligns with that of Neu & Zhivotovskiy (2020), which explored the role of abstention in the context of online prediction with expert advice.
Remark 2 (Graph Feedback).
Bandits with graph feedback (Liu et al., 2018; Alon et al., 2017) are a natural extension of the standard multi-armed bandit setting. In this framework, each arm is represented as a node in a graph, and pulling one arm reveals not only its own reward but also observations from the neighboring nodes connected to it in the graph. Our model bears a conceptual resemblance to bandits with graph feedback: it can be represented as a bipartite graph in which each stochastic arm $a$ is associated with a corresponding abstention arm $Q_{a}$. Pulling a stochastic arm $a$ reveals a reward drawn from arm $a$’s distribution, while choosing the corresponding abstention arm $Q_{a}$ yields an observation drawn from the reward distribution of the associated stochastic arm as well.
That said, there is a key structural difference between the two frameworks. In bandits with graph feedback, selecting an arm reveals not only the reward of the chosen arm but also observations from its neighboring arms. In contrast, in our model, choosing an abstention arm does not provide any information about the abstention arm itself. Instead, the mean reward (or regret) of each abstention arm is known to the learner at the beginning of the game.
@@
Our overarching goal is to design and analyze online algorithms $\pi$ that minimize their expected cumulative regrets up to and including the time horizon TT. [^3] In certain real-world applications, the time horizon $T$ may be unknown to the agent. In fact, all of our proposed methods are inherently anytime in nature, as they do not necessitate prior knowledge of the horizon. The regrets are formally defined for the two distinct settings as follows:
@@
$R^{\mathrm{RG}}_{\mu,c}(T,\pi)=\mathbb{E}\Bigg[$ $\sum_{t=1}^{T}\big((\mu_{1}-X_{t})\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\big)\Bigg].$ (1)
@@
$R^{\mathrm{RW}}_{\mu,c}(T,\pi)=T\!\cdot\!(\mu_{1}\vee c)-\mathbb{E}\Bigg[$ $\sum_{t=1}^{T}\Big(X_{t}\cdot\mathbf{1}\{B_{t}=0\}+c\cdot\mathbf{1}\{B_{t}=1\}\Big)\Bigg].$ (2)
@@
Remark 3.
It is worth mentioning that our model is a strict generalization of the canonical multi-armed bandit model (without the abstention option). Specifically, it particularizes to the canonical model as the abstention regret $c$ tends to $\infty$ in the fixed-regret setting and as the abstention reward $c$ tends to -$\infty$ in the fixed-reward setting. Nevertheless, the incorporation of an extra challenge, the abstention decision (denoted as $B_{t}$), offers the agent the potential opportunity to achieve superior performance in terms of either regret. One might wonder whether our setting reduces to a $(K+1)$-armed bandit model where the $(K+1)$-st arm models abstention. It does not; for a detailed discussion on why our framework cannot be reduced to a standard bandit problem with an additional deterministic arm, the reader is referred to Appendix B.
@@
For $x,y\in\mathbb{R}$, we denote $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. For any arm $i\in[K]$, let $N_{i}(t):=\sum_{s=1}^{t}\mathbf{1}\{A_{s}=i\}$ and $\hat{\mu}_{i}(t):=\sum_{s=1}^{t}X_{s}\mathbf{1}\{A_{s}=i\}/N_{i}(t)$ denote its total number of pulls and empirical estimate of the mean up to time $t$, respectively. In particular, we set $\hat{\mu}_{i}(t)=+\infty$ if $N_{i}(t)=0$. To count abstention records, we also use $N_{i}^{(0)}(t)$ and $N_{i}^{(1)}(t)$ to denote its number of pulls without and with abstention up to time $t$, respectively. That is, $N_{i}^{(0)}(t):=\sum_{s=1}^{t}\mathbf{1}\{A_{s}=i\text{ and }B_{s}=0\}$ and $N_{i}^{(1)}(t):=\sum_{s=1}^{t}\mathbf{1}\{A_{s}=i\text{ and }B_{s}=1\}$. Additionally, we define $\hat{\mu}_{is}$ as the empirical mean of arm $i$ based on its first $s$ pulls. Furthermore, we use $\alpha$, $\alpha_{1}$, and so forth to represent universal constants that do not depend on the problem instances (including $\mu$, $c$, $T$, $K$), with possibly different values in different contexts.
@@ 3 Fixed-Regret Setting
In terms of the arm sampling rule, our algorithm is built upon Less-Exploring Thompson Sampling (Jin et al., 2023), a minimax optimal enhancement of the celebrated Thompson Sampling (TS) algorithm (Thompson, 1933). We refer to Remark 4 for the reason behind this choice. During the initialization phase, each arm is sampled exactly once. Following that, at each time $t$, an estimated reward $a_{i}(t)$ is constructed for each arm $i\in[K]$, which is either drawn from the posterior distribution $\mathcal{N}(\hat{\mu}_{i}(t-1),1/N_{i}(t-1))$ with probability $1/K$ or set to be the empirical mean $\hat{\mu}_{i}(t-1)$ otherwise. Subsequently, the algorithm consistently pulls the arm $A_{t}$ with the highest estimated reward.
To pursue the dual optimality properties, it is essential to carefully balance the simultaneous control of asymptotic and worst-case regrets resulting from potential misjudgments in choosing the abstention option. With regard to the abstention decision rule, we propose two abstention criteria that work in tandem (as detailed in Step 5 of Algorithm 1). The first criterion is gap-dependent in nature. In particular, for each arm $i\in[K]$, we define its lower confidence bound as
@@
Then we choose to abstain if there exists an arm $i\in[K]\setminus\{A_{t}\}$ for which the difference between $\mathrm{LCB}_{i}(t)$ and the empirical mean of the arm $A_{t}$ exceeds $c$. This condition signifies that the suboptimality gap $\Delta_{A_{t}}$ is at least $c$ with high probability. The LCB in (3) is custom-tailored for our specific application, which is unique in the fact that it takes the value of the abstention regret $c$ into consideration. Notably, besides its asymptotic optimality, our choice to employ this LCB rather than other natural types of upper or lower confidence bounds in the abstention criteria (e.g., those that do not include $c$) is primarily due to the imperative to constrain the minimax regret to be $O(\sqrt{KT})$.
The second abstention criterion is gap-independent. It is motivated from the construction of worst-case scenarios as detailed in the proof of our lower bound. Under this criterion, we opt for the abstention option if $c\leq\sqrt{K/t}$, which implies that the abstention regret remains acceptably low at time $t$ in view of the worst-case scenarios.
Remark 4.
As previously highlighted, our model in the fixed-regret setting particularizes to the canonical multi-armed bandit model as the abstention regret $c$ tends to infinity. Similarly, when $c$ tends to infinity, the two abstention criteria are never satisfied, and the procedure of Algorithm 1 simplifies to that of Less-Exploring Thompson Sampling. It is worth noting that this latter algorithm is not only asymptotically optimal but also minimax optimal for the canonical model. This is precisely why we base our algorithm upon it, rather than the conventional Thompson Sampling algorithm, which has been shown not to be minimax optimal (Agrawal & Goyal, 2017).